### v1

# Leveraging high fluence and low pressure for pulsed laser deposition of high-mobility -Al2O3/SrTiO3 heterostructure growth

Abstract
High-mobility oxide heterostructures could be applied for high-frequency devices, transparent conductors, and spin-orbit logic devices. SrTiO3 is one of the most studied oxide substrate materials for heterostructures. To date, the highest SrTiO3-based charge carrier mobility at 2 K was measured in the interfacial 2-dimensional electron gas (2DEG) of -Al2O3/SrTiO3. The formation mechanism and origin of the high electron mobility are not yet fully understood. This investigation presents a successful growth protocol to synthesise high-mobility -Al2O3/SrTiO3 interfaces, and a description of the underlying growth optimisation. Furthermore, indicative features of high-mobility -Al2O3/SrTiO3, including the room-temperature sheet resistance, are presented. Signs of epitaxial and crystalline growth are found in a high-mobility sample ( cm2/Vs). Outlining the growth mechanisms and comparing 40 samples, indicates that high-fluence () and low pressure ( mbar) are essential growth parameters for high-mobility -Al2O3/SrTiO3 interfaces.
For two decades, oxide-based two-dimensional electron gasses (2DEGs) have been intensively studiedOhtomo and Hwang (2004) for their diverse features, such as superconductivity,Reyren et al. (2007) ferromagnetism,Brinkman et al. (2007) tunable Rashba spin-orbit coupling,Caviglia et al. (2010) high-mobility transport,Chen et al. (2013); Christensen et al. (2018) extreme magnetoresistance,Christensen et al. (2024) and spin-charge interconversion.Trier et al. (2020); Noël et al. (2020); Vaz et al. (2020) In recent years, several steps have been taken toward implementing such oxide interfaces. This includes the generation and detection of spin-currents in ferromagnet-free oxide nanodevicesTrier et al. (2020) and fabrication of free-standing oxide interface membranes,Sambri et al. (2020); Dahm et al. (2021); Li et al. (2022), paving the way for cointegrating oxide and silicon electronics and spintronics in everyday devices. While most studies focus on the LaAlO3/SrTiO3 interface, the highest charge carrier mobility, at 2 K, is recorded for -Al2O3/SrTiO3,Chen et al. (2013) presumably due to a band-inversion that renders the dxz/yz band dominant in electronic transport.Cao et al. (2016); Chikina et al. (2021) This attractive property allows for mean free path dependent quantum and spintronics measurements,Oltscher et al. (2014) with potential in applications including high-frequency devices,Trier, Christensen, and Pryds (2018) transparent conductors,Trier, Christensen, and Pryds (2018) and spin–orbit logic devices.Manipatruni et al. (2019)
At room temperature, longitudinal optical (LO) phonons and electron-electron scattering dominate the carrier mobility in SrTiO3.Christensen et al. (2018); Himmetoglu et al. (2014); Mikheev et al. (2015) These two scattering phenomena dominate at different carrier density ranges.
One attempt to counteract the LO phonon scattering, leveraged the non-polar, low effective electron mass, oxide BaSnO3, instead of SrTiO3, to reach room temperature mobilities on 60 , an order of magnitude higher than SrTiO3-based 2DEGs.Eom et al. (2022) The different mobilities of the two materials are attributed to the orbitals near the Fermi Energy. In SrTiO3, these orbitals are the d-band orbitals of Ti, with degenerate t2g symmetry, while it is the isotropic s-band orbitals in BaSnO3. From the less dispersive d-band orbitals, the carriers have a higher rate of inter-orbital scattering in SrTiO3 compared to s-band orbital BaSnO3.Eom et al. (2022)
At low temperatures, phonon-scattering vanishes, and the main mobility-reducing mechanism in SrTiO3 2DEGs is impurity scattering. This complicates achieving high mobility, as one of the major mechanisms forming the 2DEGs uses oxygen vacancies. Oxygen vacancies act as both dopants and scattering sites, motivating several studies into the spatial separation of oxygen vacancies and the 2DEG that arises from them.Zurhelle et al. (2020); Chikina et al. (2021); Chen et al. (2015); Huijben et al. (2013)
Here, we present an optimisation of the pulsed laser deposition (PLD) parameters of -Al2O3 grown on SrTiO3 towards forming a high-mobility 2DEG. First, the deposition chamber used in growth, along with four different protocols is presented. Subsequently, relations between low-temperature mobility and residual resistance ratio (RRR) and room-temperature sheet resistance (), respectively, are analysed. Then the interface structure of a high-mobility sample is addressed, including epitaxial strain. We then discuss the mechanisms responsible for the high-mobility interfaces throughout the growth process, from as-received substrate SrTiO3 to high-mobility -Al2O3/SrTiO3 (001) 2DEG. Based on this discussion, we compare the growth protocols and highlight indications of a lower threshold for the laser fluence and a strong pressure-dependence of the 2DEG formation.
Commercial SrTiO3(100) substratesLTD. SHINKOSHA CO. were cleaned by ultrasonication in acetone (480 s), isopropanol (IPA,480 s), and acetone:IPA 50:50 (480 s), then annealed in a 1 bar O2 atmosphere with a 60 min dwell time at 1000∘C and an up/down ramp rate of 100 ∘C/h ( 25 h in total with passive cooling). Before -Al2O3 deposition, the substrates underwent a second round of ultrasonication in acetone (480 s), IPA (480 s), and acetone:IPA 50:50 (480 s), respectively. The depositions, carried out by PLD, used a single-crystalline -Al2O3 target. The ablated plasma species forms the -Al2O3 structure on SrTiO3, as displayed in Fig. 1a). A 248 nm KrF excimer laser with a repetition rate of 1 Hz was used. We find that the detailed configuration of the growth reactor plays an important role in the optimisation process. The PLD chamber, illustrated in Fig. 1b), is directly connected to three manually controlled valves. One of the valves, V1, is connected to a mass flow controller for oxygen injection, while two other valves, V2 and V3, connect the chamber to a turbomolecular pump. This pump is further connected to a backing pump through another manual valve (not shown). V2 and V3 controlled the specific pressure when working below 10-4 mbar. V2 is a manual gate valve controlling the range of pressure, while V3 is an angle valve, allowing for fine-tuning of the pressure. V1 was open during O2 flushing and automatised proportional–integral–derivative pressure control. The thermocouple, measuring the sample temperature, resides in the heater centre, and the substrate-target distance was 4.5 0.1 cm.
Parameters | Abbreviaion | Protocol 1 | Protocol 2 | Protocol 3 | Protocol 4 |
Flushing Pressure (mbar) | 7 10-2 | NA | 7 10-2 | 7 10-2 | |
Deposition pressure (mbar) | 1-2 10-6 | 1 10-6 | 1 10-6 | 1 10-6 | |
Preablation Pressure (mbar) | |||||
Heating pressure (mbar) | 1 10-6 | 1 10-4 | Base | Base | |
Cooling pressure (mbar) | 1 10-6 | 1 10-4 | Base | 1 10-4 | |
Ramp rate (∘C/min) | 15 | 15 | 10,15 | 10,15 | |
Deposition temperature (∘C) | 650 | 550, 650 | 550 | 550, 660, 650 | |
Preablation time (min) | 10 | 10 | 10 | 10 | |
Predeposition annealing time (min) | 11,16,51 | 6-51 | 11 | 11 | |
Postdeposition annealing time (min) | 5 | 5 | 5 | 5 | |
Laser fluence (J/cm2) | 3.1-9.6 | 3.7-9.2 | 7.1-9.3 | 6.4-8.0 |
Pressure and temperature data of four protocols are displayed in Fig. 1c-f) and the different growth parameters are listed in Table 1. Protocol 1 was the first attempted protocol that yielded high-mobility samples, the other three protocols were attempts to automate and further optimise the growth process concerning reproducibility and mobility. Experiments before Protocol 1 all showed mobilities below , likely due to higher pressures in the PLD chamber.
Protocol 1, Fig. 1c), started with flushing the chamber with = 7 10-2 mbar O2 to ensure that the dominant gas in the chamber was O2. Then the chamber was evacuated to = 1-2 10-6 mbar and the target preablated for = 10 min with a shutter shielding the SrTiO3 substrate. Using , the heater, onto which the substrate was thermally and electrically attached using conducting silver paint, was heated to the deposition temperature of = 650 ∘C with a ramp rate of = 15 ∘C/min. The sample was pre-annealed for = 11 min, except where else is stated, at deposition pressure and temperature before 4 min of ablation began. After ablation, the sample was post-annealed for = 5 min at deposition pressure and temperature before the onset of the cooling. The cooling was, like the heating, set to a ramp rate of = 15 ∘C/min, below 300 ∘C the ramp rate decreased since the sole cooling power came from the local, low-pressure, environment.
Protocol 2, as displayed in Fig. 1d), differed from Protocol 1 mainly by = = 1 10-4 mbar O2. For this protocol, different durations were tested and initial flushing with O2 was excluded.
In Protocol 3, shown in Fig. 1e), = 550 ∘C and V2 and V3 left completely open without any O2 in-flow during the whole process except for the initial O2 flush. This decreases and to base pressure, which depend on the heating and pumping.
Protocol 4 was identical to Protocol 3 with the exception that = 10-4 mbar, as displayed in Fig. 1f).
The -Al2O3 growth thicknesses were measured by X-ray reflectivity (XRR) using a film grown by Protocol 1 (-Al2O3 8.5 nm). The growth time (4 min) was the same for all samples. X-ray diffraction (XRD), XRR, and atomic force microscopy (AFM), were used to assess the crystal structure.
Electrical connections were made by bonding with Al ultrasonic wire-bonding, in van der Pauw geometry. Electronic transport below room temperature was measured using a Cryogen-Free Measurement System equipped with a 16 T magnet.
For the samples of this study, explicit low-temperature mobility measurements need cooling and testing in vdP and Hall configurations with magnetic field sweeps. Indicative room temperature parameters allow quicker mobility optimisation. As displayed in Fig. 2a) such indicative features are extractable from the sheet resistance temperature-dependency. Both high (4.2 at 2 K) and low (6 at 2 K) mobility samples here show metallic behaviour but differ in the absolute resistance and residual resistance ratio (RRR(10 K) = Rs(300 K)/Rs(10 K)). The mobility as a function of carrier density is shown in Fig. 2b). A peak in mobility was reportedChristensen et al. (2018) at 4 1014 cm-2, within a range of carrier densities not accessed in this investigation. Plotting the mobility as a function of room temperature sheet resistance (), displayed in Fig. 2c), reveals a qualitative mirror image of Fig. 2b). The inset indicates the log-log linear dependency between the carrier density and , explaining these similarities. The dependency between mobility and , through the carrier density, reveals as a relevant indicative value to estimate the low-temperature mobility.
One can define a relevant upper threshold for high mobility in at 10 k/sq. A lower threshold is not determinable from Fig. 2, although some level of conductance between the top and bottom of a sample can provide an alternative, as 3D conductance is undesirable for 2D features. A threshold 3D conductance requires additional investigations.
The total carrier density is often temperature independent, hence RRR can be understood as an indirect measurement of . Figure 2d) exhibits the mobility at 10 K as a function of RRR(10 K). RRR(10 K) increases with increasing mobility, yielding a range for high-mobility ( 104) samples at RRR(10 K) 3 103. This can be observed because both and are proportional to the approximately temperature-independent .
To gain insights into the relations between the crystalline quality and the mobility, sample 9 ( ) was measured using AFM, XRR, and XRD. The visible terraces in AFM after deposition, as displayed in Fig. 3a), indicate crystalline and epitaxial growth of the -Al2O3 thin film.
To estimate the thin film thicknesses and confirm epitaxial crystalline growth XRD and XRR were measured. In Fig. 3b) the XRR data is displayed along with a fit determining the -Al2O3 thickness to nm. Fig. 3c) shows XRD 2 scan of the -Al2O3 sample 9 grown on SrTiO3 (001). -Al2O3 grows with a positive out-of-plane strain of 0.721(1)%. Calculated by using the c-lattice parameter of bulk -Al2O3 (7.911 Å).Rongsheng and Snyder (1991)
A thorough understanding of the 2DEG formation mechanism is essential to optimise the mobility in -Al2O3/SrTiO3 further. As previously described in ref. Hvid-Olsen et al. (2024), annealing in oxygen causes a transition from Rs(RT) = 310 /sq to Rs(RT) 950 M/sq. This points towards oxygen vacancies being the primary source of the 2DEG formation.
Figure 4a) illustrates how the samples go through six regimes in the growth process. Regime I involves annealing at atmospheric pressure and high temperatures. The sample is then loaded into the PLD chamber, initiating Regime II during heating under low pressure. Regime III occurs as the SrTiO3 surface remains exposed to low pressure at the growth temperature. Deposition begins in Regime IV and ends before Regime V, which includes cooldown. Regime VI covers subsequent sample storage.
Regime I is illustrated in Fig. 4b), where the arrow indicates the direction and magnitude of the oxygen atom diffusion to restore thermodynamic equilibrium between the SrTiO3 lattice and the environment.
Since the temperature reaches 1000 oC in O2 environment, the oxygen atoms become highly mobile and vacancies are filled. This creates a reproducible starting point regarding oxygen distribution. During the subsequent cooling, the atomic mobility decreases, making the oxygen occupation in the sample reflect the environment, functioning as a pressure memory of the furnace equilibrium.
In regimes II and III, illustrated in Fig. 4c), the substrate is loaded in the PLD chamber, where the pressure is adjusted to 10-4 mbar. Now, the pressure balance is reversed. With the low-pressure environment and the memory of the furnace pressure in the lattice, the oxygen atoms will diffuse out of the lattice to restore thermodynamic equilibrium. The topmost atomic layer will experience the fastest reduction since this layer is directly exposed to the environment.
Once a vacancy is generated in the topmost layer, the system has moved slightly towards equilibrium. This decreases the oxygen diffusion, as illustrated in Fig. 4d) and allows the oxygen vacancies to move further into the bulk SrTiO3,. When the substrate reaches deposition temperature, a laser pulse is sent into the target, entering regime IV. The laser pulse is ablated into a high-energy plasma. When the plasma reaches the substrate a bombardment effect occurs, as illustrated in Fig. 4e). The plasma adds energy on the scale of tens of eVSambri et al. (2012); von Soosten et al. (2019) to the top layers of the lattice, enabling the escape of light elements. In SrTiO3, the light element is oxygen.
Because -Al2O3 only possess a single cationic element, one can increase the fluence without considering the effect on the cationic stoichiometry. Hence, one can utilise the fluence to access a larger range of energetic minima associated with crystal growth of -Al2O3 compared to e.g. LaAlO3 or LaTiO3.
As the -Al2O3 thin film is grown in the oxygen-deficient environment, it further reduces SrTiO3 since the oxygen affinity of Al is higher than that of Ti and Sr. The combination of the high temperature and the kinetic energy from the plasma ensures high atomic mobility.Chen et al. (2011); Christensen et al. (2017); von Soosten et al. (2019) This corresponds to regime IV and V and is illustrated in Fig. 4f).
At last, during sample cooldown (regime V), for Protocol 2 and 4, the pressure is increased from deposition pressure to 10-4 mbar. This may cause oxygen recombination, as displayed in Fig. 4g), where an atmospheric oxygen molecule is split up and added to the lattice, annihilating two vacancies.Christensen et al. (2017); von Soosten et al. (2019); Hvid-Olsen et al. (2022) This effect may also occur during subsequent storage corresponding to regime VI.
Samples 1, 2, 3, 9, and 10 grown by Protocol 1, as already displayed in Fig. 2, were confirmed to host high-mobility 2DEGs. Protocol 1, however, has some weaknesses. Stabilising a constant low pressure during the temperature profile is challenging, often failing to reach the intended low pressure during heating, probably due to outgassing from the silver paint and heater. This is likely the reason for the large fluctuation in , as shown in Supplementary Material Table S1. Additionally, manual pressure control is time-consuming (5 hours) and less reproducible than automated control.
Using Protocol 1, only samples grown with = 6 displayed high mobility. This points towards a lower threshold of the fluence consistent with a high kinetic energy, allowing access to a greater range of energy minima of the plasma species. It is worth to note that the fluences reported in ref. 5 and 6, were 1.5 J/cm2 and 3.8 J/cm2, respectively. Hence, one should regard the high-fluence threshold of this investigation as a trend, but with chamber-specific deviations. Such deviation is reasonable, since the optics from chamber to chamber can vary substantially.
To increase the reproducibility and time efficiency of the growth, Protocol 2, with mbar, was attempted to create samples 19-33. In contrast to Protocol 1 however, Protocol 2 shows measurable conductance R2G(RT) at room temperature between the chip carrier back plane and any of the four corners on the upper surface of the sample. Thus, Protocol 2 yields bulk conductance before high mobility 2DEGs, when decreasing the sheet resistance by extending .
The tendency to create 3D conductance before high-mobility 2DEGs, could be explained by the balance between temperature and pressure as driving forces for oxygen vacancy formation. During regimes I-III when SrTiO3 is uncapped by the thin film, the oxygen occupancy in the area where the interface will be created can be engineered. Leveraging high temperature as a driving force, one can expect a homogeneous oxygen vacancy distribution across the substrate thickness, since the temperature, from the heater on the backside of the substrate, is lower at the substrate front than further into the substrate. If low pressure, in contrast, is the driving force, the surface of the substrate will be significantly more affected than the bulk substrate, because of its direct exposure. However, elevated atomic mobilities are a necessity, making room-temperature depositions unfavourable. Nonetheless, with the lower temperatures during heating, oxygen vacancy formation in the surface layers rather than the bulk could occur if the pressure is low enough to dominate the formation. This highlights the importance of low pressure during heating, in achieving oxygen vacancy-driven 2DEG formation.
Attempting to leverage this mechanism for the formation of surface oxygen vacancies, Protocol 3 was applied for growing samples 34-36, as shown in Table S1. This protocol also yields 3D conductance before high mobility 2DEG formation. In this case, however, the 3D conductance likely develops during post-annealing and cooldown when the pressure is at its lowest, the temperature is still significant, and the potential 2DEG area is shielded by the -Al2O3, acting as a capping layer that keeps the oxygen from escaping from the 2DEG area.
To counteract both the lack of surface oxygen vacancies from Protocol 2 and the increase in bulk oxygen vacancies in Protocol 3, Protocol 4 was developed as displayed by samples 37-40. Here, the bulk conductance does not reach measurable ranges, but the 2DEG resistance tends to increase with increasing growth temperature, rendering the protocol unsuited for high-mobility 2DEG growth. This can be attributed to oxygen recombination while the pressure is 10-4 mbar.
Within this investigation, four different growth protocols attempting to grow -Al2O3 on SrTiO3 are presented alongside small variations within each of them. The purpose was to create a 2DEG with high carrier mobility. This purpose was initially reached by Protocol 1, displaying a growth window of such epitaxial, crystalline, high mobility -Al2O3/SrTiO3 2DEGs. Protocol 1 however, is highly dependent on manual work, decreasing the reproducibility. Hence, the subsequent protocols attempted to stay within the high mobility growth window while automating the work. To optimise that process, the residual resistance ratio (RRR) and the room-temperature sheet resistance (Rs) were identified as indicative parameters for estimating the mobility at 10 K without time-consuming low-temperature measurements between every growth. For both RRR and Rs, the reason for the indicative implications is identified as the approximate temperature independence of the carrier concentration of the -Al2O3/SrTiO3 2DEGs.
Four different mechanisms for oxygen-vacancy-driven 2DEG formation were discussed. These were SrTiO3-environment equilibrium, bombardment, reduction of SrTiO3 by -Al2O3, and oxygen recombination. Considering these mechanisms and comparing 40 samples grown with the four different protocols, pointed toward a low limit for high-mobility -Al2O3/SrTiO3 growth of the fluence at 3 J/cm2, and a necessity of using low-pressure to induce oxygen vacancies in the surface of SrTiO3 before deposition.
We thank R. T. Dahm, C. N. Lobato, V. Rosendal, and C. E. N. Petersen for helpful discussions.
T.H.O., A. C. and F.T. acknowledge support by research grant 37338 (SANSIT) from Villum Fonden. T.S.J. acknowledges support from the Novo Nordic Foundation Challenge Program, grant no. NNF21OC0066526 (BioMag).